The probability of John failing modern studies (Event M) is 0.26, while the probability that he fails environment science (Event E) is 0.35. The probability that he fails both subjects is 0.18. What is the probability that he passes both the subjects?


1
Expert's answer
2021-11-22T06:15:56-0500
"P(M\cup E)=P(M)+P(E)-P(M\cap E)"

"=0.26+0.35-0.18=0.43"

"P(passes\ both)=1-P(M\cup E)"

"=1-0.43=0.57"